2. The equation D=1.2 square root of h. gives the distance, D, in miles that a person can see to the horizon from a height, h, in feet.
a. Solve this equation for h.
b. Long’s Peak in the Rocky Mountain National Park, is 14,255 feet in elevation. How far can you see to the horizon from the top of Long’s Peak? Can you see Cheyenne, Wyoming (about 89 miles away)? Explain your answer.

Answer by checkley77(12569)

D=1.2SQRTH
SQRTH=D/1.2 SQUAERE BOTH SIDES.
H=(D/1.2)^2
H=D^2/1.44 ANSWER.
14,255=D^2/1.44
D^2=14,525*1.44
D^2=20,916
D=144.62 MILES CAN BE SEEN.
YES. CHEYANNE CAN BE SEEN BECAUSE 89 MILES IS LESS THAN 144.62 MILES.